# The Low-Temperature Phenomenology of Gap Inhomogeneity in the Cuprate Superconductors: High-Energy Granularity, Low-Energy Homogeneity, and Spectral Kinks

### v1

@@ VIII.3 Near-Node Gap Measurements
Figure 26: Left: Plot of the extracted inverse nodal slope ($\delta_{N}$, $y$-axis) and the antinodal gap ($\Delta_{0}$, $x$-axis) from $dI/dV$ spectra taken from Bi-2212 at various dopings. From Ref. Pushp et al. (2009), Pushp et al. Reprinted with permission from AAAS. Right: Plots of the binned $\Delta_{L}$ vs. $\Delta_{S}$ with $\Delta=0.08$, $l=16$ (purple); $\Delta=0.16$, $l=8$ (blue); and $\Delta=0.32$, $l=4$ (orange), taken from simulated data. The binning is done by $\Delta_{S}$, and corresponds to data previously shown in Fig. 22. The dashed line indicates $\Delta_{T}$ = $\Delta_{S}$. The size of each marker is proportional to the number of spectra within the corresponding bin. All three of these scenarios fall within the $l\approx\xi$ regime.
Pushp et al. perform a systematic doping-dependent analysis and find that the underdoped cuprates exhibit a “universal” low-energy regime where “nodal gap”—i.e., an estimate of the order parameter derived from either the slope of the local density of states near $\omega=0$ (which is equivalent to our definition of the low-energy gap $\Delta_{L}$) or from the result of their fitting procedure taken at the gap node—is strikingly doping-independent from the strongly underdoped regime all the way to optimal doping Pushp et al. (2009). This universality is lost beyond optimal doping: for overdoped cuprates, they find that nodal gap decreases with increasing doping. They also find that as doping is lowered, the kink energies become smaller and that $\Delta_{L}$ is less than all binned $\Delta_{S}$ for heavily underdoped Bi-2212 ($T_{c}=35$ K and $T_{c}=58$ K). It should be noted that in this regime, the large-$\Delta_{S}$ spectra show extremely rounded coherence peaks that do not resemble those arising from a mean-field model of $d$-wave superconductivity; this is also seen in the STS data from McElroy et al. taken in this regime McElroy et al. (2005a).
@@
Moderately underdoped and optimally doped Bi-2212 on the other hand appears to be well described by a doping-independent $\overline{\Delta}$, which is necessary to ensure the observed behavior of $\Delta_{L}$. The fact that the kink energy as a fraction of the average spectral gap increases with increasing doping is consistent with the decrease of the inhomogeneity length scale $l$ (with fixed $\xi$, since $\overline{\Delta}$ is roughly constant) and/or the decrease of the width of the inhomogeneity distribution as doping is increased. It should be noted that McElroy et al. McElroy et al. (2005a), Gomes et al. Gomes et al. (2007), Alldredge et al. Alldredge et al. (2008), and Pushp et al. Pushp et al. (2009) find that the width of the $\Delta_{S}$ distribution decreases with increasing doping. Both possibilities (decreasing $l$ or width of the $\Delta_{T}$ distribution) are consistent with this observation. Recall that in Fig. 22, we see that even though the underlying order-parameter distributions have the same mean value $\overline{\Delta}$, the resulting average of the spectral gaps can nevertheless vary as a function of $l$, with the maximum difference between $\overline{\Delta}$ and the average $\Delta_{S}$ occuring when $l/\xi\approx 2$-$4$. Taking this into consideration, the average value of $\Delta_{S}$ as inferred from experiment is likely to be an overestimate of $\overline{\Delta}$ in the cuprates. The observed increase of the average $\Delta_{S}$ and the accompanying decrease in $\Delta_{K}$ with decreasing doping could likely arise from an increase in the intrinsic inhomogeneity length scale ratio $l$. Much of the doping-dependent phenomenology of $\Delta_{S}$, $\Delta_{L}$ and $\Delta_{K}$ described in Ref. Pushp et al. (2009), in Fig. 6b of Ref. Alldredge et al. (2008), and in Fig. 13b in Ref. Alldredge et al. (2013) can thus be captured by the plots in Fig. 22 which show what happens to these various energy scales when both $\overline{\Delta}$ and the inhomogeneity width are both kept fixed while $l$ increases (keeping $\xi$ fixed). Having said this, we note that these behaviors could just as well be described by a decrease in the width of the inhomogeneity distribution with increasing doping: the kink energy is determined by the smallest spectral gaps which would increase when the width of the inhomogeneity is decreased while holding $\overline{\Delta}$ constant.
@@
We note finally that Fig. 12c, which describes the joint distribution of $\Delta_{S}$ and $\Delta_{L}$ for the $l\approx\xi$ case, is very similar to Fig. 4a of Pushp et al. Pushp et al. (2009) (which describes the same quantities, but using bin-averages) for a variety of dopings ranging from moderate underdoping ($T_{c}=58$ K) to moderate overdoping ($T_{c}=65$ K). (An explicit comparison of the experimental data from Pushp et al. with our $l\approx\xi$ simulated results with $\overline{\Delta}=0.08$, $\overline{\Delta}=0.16$, and $\overline{\Delta}=0.32$, using bin-averaged data to replicate fully the presentation of the Pushp et al. results, is shown in Fig. 26.) Both the experimental and simulated data show similar features: the $\Delta_{L}$ distributions are much narrower than those of $\Delta_{S}$, and, rather strikingly, the joint distribution mainly extends out towards the $\Delta_{L}<\Delta_{S}$ portion of the plot from $\Delta_{L}\approx\Delta_{S}$, with comparatively little weight in the $\Delta_{L}>\Delta_{S}$ portion of the plot. Within our model, this form of the joint distribution is due to the fact that a relatively large minimum spectral gap is present even in regions where the order parameter is zero, so much of the weight has piled onto the smallest $\Delta_{S}$’s (where this weight would have been more spread out in the case of the joint distribution of $\Delta_{T}$ and $\Delta_{L}$). That said, we note that in Fig. 26, the experimental joint distributions of $\Delta_{S}$ and $\Delta_{L}$ are such that these lie entirely within the $\Delta_{L}<\Delta_{S}$ portion of the plot. We are not sure as to why this is the case. In our numerical work, in the $l\approx\xi$ regime, we generally find that the average spectral gap $\Delta_{S}$ is generally larger than the average low-energy gap $\Delta_{L}$, but nevertheless a small but nonzero fraction of spectra satisfy $\Delta_{L}>\Delta_{S}$. As far as the cuprates are concerned, this shift may signal the presence of subtle self-energy effects not included in our analysis that renormalize the spectral gap upward at higher energies, but should these be the reason underlying this effect, it is remarkable that that these persist even at overdoping, where correlation effects are expected to be weaker. It should be said that this shift is less pronounced when $\overline{\Delta}$ is small, so correlation effects may not be needed to be invoked in the overdoped case; see the plot for $\overline{\Delta}=0.08$ in Fig. 26, which shows that nearly all of the weight of the joint distribution lies below the $\Delta_{L}=\Delta_{S}$ line.
@@ VIII.4 Self-Energy Fits
Many analyses of STS data on the cuprates employ a procedure in which the individual $dI/dV$ spectra are fitted to a “Dynes-like” formula involving a $d$-wave order parameter, a suitable normal-state tight-binding model, and a scattering rate $\Gamma$ to account for the broadening of spectra. Various models of scattering rates have been used in the literature. Commonly used is an energy-independent but spatially varying scattering rate Pasupathy et al. (2008); Tromp et al. (2023), which ensures that the broadening is uniform at all energies. More sophisticated approaches include an energy- and position-dependent scattering rate (e.g., $\Gamma(\mathbf{r}_{i},\omega)\propto\omega$, as used in Refs. Alldredge et al. (2008) and Alldredge et al. (2013)) which takes into account inelastic scattering from electron-electron interactions. Such an energy-dependent model is able to account on a phenomenological level for the shorter and broadened coherence peaks in large-gap spectra taken from underdoped cuprates.
@@
It is interesting to revisit these scattering-rate estimates in light of our studies of $d$-wave superconductors with inhomogeneous pairing in the $l\approx\xi$ regime. We generally find that without the explicit inclusion of an energy-dependent scattering rate (recall that we use an energy-independent $\Gamma$), the coherence peaks of large-$\Delta_{S}$ regions are already shorter than expected from a clean bulk $d$-wave superconductor with the same $\Delta_{S}$ and $\Gamma$. These suggest that the scattering-rate estimates made in Refs. Alldredge et al. (2008) and Alldredge et al. (2013) are an overestimate, as their fits do not take into account the inhomogeneity of the order parameter. The fitting procedure they employ assumes that each local spectrum is captured by a clean model of a $d$-wave superconductor with some $\Delta_{T}(\mathbf{r}_{i})$ and $\Gamma(\mathbf{r}_{i},\omega)$. However, we have already seen from our examples earlier in this paper that the identification of the local $\Delta_{S}$ and the true local $\Delta_{T}$ is not exact, especially in the $l\approx\xi$ case relevant to Bi-2212 and Bi-2201. As mentioned earlier, the origin of the shorter coherence peaks in the $l\approx\xi$ regime are the low-energy homogeneity and the kinks: as the low-energy spectra are homogeneous, the large-gap spectra feature much more spectral weight at low energies compared to the bulk $d$-wave superconductor with the same large spectral gap, leaving less spectral weight for the high-energy coherence peak and rendering it comparatively shorter. While the large-gap spectra do feature considerable broadening to suggest that inelastic scattering plays an important role (especially in heavily underdoped cuprates), the amount of inelastic scattering needed to reproduce the spectra appears to be less than originally estimated.
@@
The scattering-rate estimates made in Refs. Alldredge et al. (2008) and Alldredge et al. (2013) can be reinterpreted in light of our numerics in the following way. These quantify the amount of spectral weight that is transferred away from the coherence peaks to other parts of the spectrum, regardless of the mechanism for the spectral-weight transfer (inhomogeneity, inelastic scattering, etc.). Order-parameter inhomogeneity transfers spectral weight to the kink, while inelastic scattering broadens out the coherence peak directly and redistributes spectral weight mainly in the vicinity of the peak. These scattering-rate estimates do not distinguish between these two contributions to spectral-weight transfer. In principle, it would be possible to disentangle these two effects by numerically computing the spectra of the inhomogeneous system and iteratively adjusting $\Delta_{T}(\mathbf{r}_{i})$ and $\Gamma(\mathbf{r}_{i},\omega)$ until decent fits to the experimental spectra are obtained.
@@ VIII.5 Extinction of Quasiparticle Scattering Interference
Quasiparticle scattering interference is a ubiquitous phenomenon in the cuprates; however, it is famously known to disappear in the underdoped cuprates when the bias voltage is raised above a critical threshold. This phenomenon (dubbed “QPI extinction”) has been observed in underdoped and optimally doped Bi-2212 Kohsaka et al. (2008); Fujita et al. (2014) and Bi-2201 He et al. (2014); Webb et al. (2019), but not at overdoping, where QPI continues to be present at all energies below the mean spectral gap of the system. Strikingly, Kohsaka et al. find that the energy at which QPI disappears is the same energy at which the homogeneity of the low-energy states is lost Kohsaka et al. (2008). This last observation provides an important clue to pinpointing the true origin of QPI extinction.
@@
Our results on inhomogeneous $d$-wave superconductors provide a possible reason for this phenomenon. At low energies below $\Delta_{K}$, the quasiparticle states are homogeneous, so coherent scattering can occur when these quasiparticles encounter short-ranged quenched disorder. In this regime, the QPI response should be that of a uniform $d$-wave superconductor with order parameter $\overline{\Delta}$. Above $\Delta_{K}$, on the other hand, the states are now inhomogeneous: the LDOS now reflects the local value of $\Delta_{T}$, and sharp differences between the LDOS on nearby sites are present. We had seen in our discussion of Figs. 10 that for energies $\omega>\Delta_{K}$, the system divides itself into three classes of regions corresponding to whether the local spectral gap $\Delta_{S}$ at any particular site satisfies $\omega\approx\Delta_{S}$, $\omega>\Delta_{S}$, or $\omega<\Delta_{S}$, with the quasiparticle states localized within any one of these three classes of regions. Thus, the coherent propagation of wavelike quasiparticle excitations across the entire system ceases to be possible when $\omega>\Delta_{K}$, and therefore QPI is no longer present. Details of this argument, making use of a model of inhomogeneous $d$-wave superconductivity with weak scatterers present, will be presented in a future manuscript Sulangi .
@@
While compelling, a crucial weakness of this picture is that it does not explain why $\Delta_{K}$ (i.e., the energy where both low-energy homogeneity and QPI disappear) exactly coincides with the energy where the tips of the contours of constant energy (CCE) of the Bogoliubov quasiparticles crosses the antiferromagnetic zone boundary. Within our model of inhomogeneous $d$-wave superconductivity, there is no reason why $\Delta_{K}$ should match this energy, which would be set by the band-structure parameters and $\overline{\Delta}$. This additional requirement likely signals that additional ingredients beyond an inhomogeneous $d$-wave order parameter should be present; this is likely where the broken-symmetry states at high energies play a key role. Other proposals for resolving the problem of QPI extinction have been put forth, invoking coexisting antiferromagnetic order Andersen and Hirschfeld (2009) and modulated hoppings near impurities Vishik et al. (2009); these proposals explain why QPI disappears when the tips of the CCEs extend past the AF zone boundary, but not why this energy coincides with the loss of homogeneity. A full explanation of this phenomenon will likely have to involve a combination of inhomogeneity and some of these other ingredients, in addition to squaring fully with the highly correlated nature of the underdoped superconducting state.

### v2

@@ VIII.3 Near-Node Gap Measurements
Figure 25: Left: Plot of the extracted inverse nodal slope ($\delta_{N}$, $y$-axis) and the antinodal gap ($\Delta_{0}$, $x$-axis) from $dI/dV$ spectra taken from Bi-2212 at various dopings. From Ref. [12], Pushp et al. Reprinted with permission from AAAS. Right: Plots of the binned $\Delta_{L}$ vs. $\Delta_{S}$ with $\Delta=0.08$, $l=16$ (purple); $\Delta=0.16$, $l=8$ (blue); and $\Delta=0.32$, $l=4$ (orange), taken from simulated data. The binning is done by $\Delta_{S}$, and corresponds to data previously shown in Fig. 21. The dashed line indicates $\Delta_{T}$ = $\Delta_{S}$. The size of each marker is proportional to the number of spectra within the corresponding bin. All three of these scenarios fall within the $l\approx\xi$ regime.
Pushp et al. perform a systematic doping-dependent analysis and find that the underdoped cuprates exhibit a “universal” low-energy regime where “nodal gap”—i.e., an estimate of the order parameter derived from either the slope of the local density of states near $\omega=0$ (which is equivalent to our definition of the low-energy gap $\Delta_{L}$) or from the result of their fitting procedure taken at the gap node—is strikingly doping-independent from the strongly underdoped regime all the way to optimal doping [12]. This universality is lost beyond optimal doping: for overdoped cuprates, they find that nodal gap decreases with increasing doping. They also find that as doping is lowered, the kink energies become smaller and that $\Delta_{L}$ is less than all binned $\Delta_{S}$ for heavily underdoped Bi-2212 ($T_{c}=35$ K and $T_{c}=58$ K). It should be noted that in this regime, the large-$\Delta_{S}$ spectra show extremely rounded coherence peaks that do not resemble those arising from a mean-field model of $d$-wave superconductivity; this is also seen in the STS data from McElroy et al. taken in this regime [6].
@@
Moderately underdoped and optimally doped Bi-2212 on the other hand appears to be well described by a doping-independent $\overline{\Delta}$, which is necessary to ensure the observed behavior of $\Delta_{L}$. The fact that the kink energy as a fraction of the average spectral gap increases with increasing doping is consistent with the decrease of the inhomogeneity length scale $l$ (with fixed $\xi$, since $\overline{\Delta}$ is roughly constant) and/or the decrease of the width of the inhomogeneity distribution as doping is increased. It should be noted that McElroy et al. [6], Gomes et al. [9], Alldredge et al. [10], and Pushp et al. [12] find that the width of the $\Delta_{S}$ distribution decreases with increasing doping. Both possibilities (decreasing $l$ or width of the $\Delta_{T}$ distribution) are consistent with this observation. Recall that in Fig. 21, we see that even though the underlying order-parameter distributions have the same mean value $\overline{\Delta}$, the resulting average of the spectral gaps can nevertheless vary as a function of $l$, with the maximum difference between $\overline{\Delta}$ and the average $\Delta_{S}$ occuring when $l/\xi\approx 2$-44. Taking this into consideration, the average value of $\Delta_{S}$ as inferred from experiment is likely to be an overestimate of $\overline{\Delta}$ in the cuprates. The observed increase of the average $\Delta_{S}$ and the accompanying decrease in $\Delta_{K}$ with decreasing doping could likely arise from an increase in the intrinsic inhomogeneity length scale ratio $l$. Much of the doping-dependent phenomenology of $\Delta_{S}$, $\Delta_{L}$ and $\Delta_{K}$ described in Ref. [12], in Fig. 6b of Ref. [10], and in Fig. 13b in Ref. [14] can thus be captured by the plots in Fig. 21 which show what happens to these various energy scales when both $\overline{\Delta}$ and the inhomogeneity width are both kept fixed while $l$ increases (keeping $\xi$ fixed). Having said this, we note that these behaviors could just as well be described by a decrease in the width of the inhomogeneity distribution with increasing doping: the kink energy is determined by the smallest spectral gaps which would increase when the width of the inhomogeneity is decreased while holding $\overline{\Delta}$ constant. This is corroborated in Appendix A.
@@
We note finally that Fig. 11c, which describes the joint distribution of $\Delta_{S}$ and $\Delta_{L}$ for the $l\approx\xi$ case, is very similar to Fig. 4a of Pushp et al. [12] (which describes the same quantities, but using bin-averages) for a variety of dopings ranging from moderate underdoping ($T_{c}=58$ K) to moderate overdoping ($T_{c}=65$ K). (An explicit comparison of the experimental data from Pushp et al. with our $l\approx\xi$ simulated results with $\overline{\Delta}=0.08$, $\overline{\Delta}=0.16$, and $\overline{\Delta}=0.32$, using bin-averaged data to replicate fully the presentation of the Pushp et al. results, is shown in Fig. 25.) Both the experimental and simulated data show similar features: the $\Delta_{L}$ distributions are much narrower than those of $\Delta_{S}$, and, rather strikingly, the joint distribution mainly extends out towards the $\Delta_{L}<\Delta_{S}$ portion of the plot from $\Delta_{L}\approx\Delta_{S}$, with comparatively little weight in the $\Delta_{L}>\Delta_{S}$ portion of the plot. Within our model, this form of the joint distribution is due to the fact that a relatively large minimum spectral gap is present even in regions where the order parameter is zero, so much of the weight has piled onto the smallest $\Delta_{S}$’s (where this weight would have been more spread out in the case of the joint distribution of $\Delta_{T}$ and $\Delta_{L}$). That said, we note that in Fig. 25, the experimental joint distributions of $\Delta_{S}$ and $\Delta_{L}$ are such that these lie entirely within the $\Delta_{L}<\Delta_{S}$ portion of the plot. We are not sure as to why this is the case. In our numerical work, in the $l\approx\xi$ regime, we generally find that the average spectral gap $\Delta_{S}$ is generally larger than the average low-energy gap $\Delta_{L}$, but nevertheless a small but nonzero fraction of spectra satisfy $\Delta_{L}>\Delta_{S}$. As far as the cuprates are concerned, this shift may signal the presence of subtle self-energy effects not included in our analysis that renormalize the spectral gap upward at higher energies, but should these be the reason underlying this effect, it is remarkable that that these persist even at overdoping, where correlation effects are expected to be weaker. It should be said that this shift is less pronounced when $\overline{\Delta}$ is small, so correlation effects may not be needed to be invoked in the overdoped case; see the plot for $\overline{\Delta}=0.08$ in Fig. 25, which shows that nearly all of the weight of the joint distribution lies below the $\Delta_{L}=\Delta_{S}$ line.
@@ VIII.4 Self-Energy Fits
Many analyses of STS data on the cuprates employ a procedure in which the individual $dI/dV$ spectra are fitted to a “Dynes-like” formula involving a $d$-wave order parameter, a suitable normal-state tight-binding model, and a scattering rate $\Gamma$ to account for the broadening of spectra. Various models of scattering rates have been used in the literature. Commonly used is an energy-independent but spatially varying scattering rate [11, 16], which ensures that the broadening is uniform at all energies. More sophisticated approaches include an energy- and position-dependent scattering rate (e.g., $\Gamma(\mathbf{r}_{i},\omega)\propto\omega$, as used in Refs. [10] and [14]) which takes into account inelastic scattering from electron-electron interactions. Such an energy-dependent model is able to account on a phenomenological level for the shorter and broadened coherence peaks in large-gap spectra taken from underdoped cuprates.
@@
It is interesting to revisit these scattering-rate estimates in light of our studies of $d$-wave superconductors with inhomogeneous pairing in the $l\approx\xi$ regime. We generally find that without the explicit inclusion of an energy-dependent scattering rate (recall that we use an energy-independent $\Gamma$), the coherence peaks of large-$\Delta_{S}$ regions are already shorter than expected from a clean bulk $d$-wave superconductor with the same $\Delta_{S}$ and $\Gamma$. These suggest that the scattering-rate estimates made in Refs. [10] and [14] are an overestimate, as their fits do not take into account the inhomogeneity of the order parameter. The fitting procedure they employ assumes that each local spectrum is captured by a clean model of a $d$-wave superconductor with some $\Delta_{T}(\mathbf{r}_{i})$ and $\Gamma(\mathbf{r}_{i},\omega)$. However, we have already seen from our examples earlier in this paper that the identification of the local $\Delta_{S}$ and the true local $\Delta_{T}$ is not exact, especially in the $l\approx\xi$ case relevant to Bi-2212 and Bi-2201. As mentioned earlier, the origin of the shorter coherence peaks in the $l\approx\xi$ regime are the low-energy homogeneity and the kinks: as the low-energy spectra are homogeneous, the large-gap spectra feature much more spectral weight at low energies compared with the bulk $d$-wave superconductor with the same large spectral gap, leaving less spectral weight for the high-energy coherence peak and rendering it comparatively shorter. While the large-gap spectra do feature considerable broadening to suggest that inelastic scattering plays an important role (especially in heavily underdoped cuprates), the amount of inelastic scattering needed to reproduce the spectra appears to be less than originally estimated.
@@
The scattering-rate estimates made in Refs. [10] and [14] can be reinterpreted in light of our numerics in the following way. These quantify the amount of spectral weight that is transferred away from the coherence peaks to other parts of the spectrum, regardless of the mechanism for the spectral-weight transfer (inhomogeneity, inelastic scattering, etc.). Order-parameter inhomogeneity transfers spectral weight to the kink, while inelastic scattering broadens out the coherence peak directly and redistributes spectral weight mainly in the vicinity of the peak. These scattering-rate estimates do not distinguish between these two contributions to spectral-weight transfer. In principle, it would be possible to disentangle these two effects by numerically computing the spectra of the inhomogeneous system and iteratively adjusting $\Delta_{T}(\mathbf{r}_{i})$ and $\Gamma(\mathbf{r}_{i},\omega)$ until decent fits to the experimental spectra are obtained.
@@ VIII.5 Extinction of Quasiparticle Scattering Interference
Quasiparticle scattering interference is a ubiquitous phenomenon in the cuprates; however, it is famously known to disappear in the underdoped cuprates when the bias voltage is raised above a critical threshold. This phenomenon (dubbed “QPI extinction”) has been observed in underdoped and optimally doped Bi-2212 [28, 29] and Bi-2201 [30, 31], but not at overdoping, where QPI continues to be present at all energies below the mean spectral gap of the system. Strikingly, Kohsaka et al. find that the energy at which QPI disappears is the same energy at which the homogeneity of the low-energy states is lost [28]. This last observation provides an important clue to pinpointing the true origin of QPI extinction.
@@
Our results on inhomogeneous $d$-wave superconductors provide a possible reason for this phenomenon. At low energies below $\Delta_{K}$, the quasiparticle states are homogeneous, so coherent scattering can occur when these quasiparticles encounter short-ranged quenched disorder. In this regime, the QPI response should be that of a uniform $d$-wave superconductor with order parameter $\overline{\Delta}$. Above $\Delta_{K}$, on the other hand, the states are now inhomogeneous: the LDOS now reflects the local value of $\Delta_{T}$, and sharp differences between the LDOS on nearby sites are present. We had seen in our discussion of Figs. 9 that for energies $\omega>\Delta_{K}$, the system divides itself into three classes of regions corresponding to whether the local spectral gap $\Delta_{S}$ at any particular site satisfies $\omega\approx\Delta_{S}$, $\omega>\Delta_{S}$, or $\omega<\Delta_{S}$, with the quasiparticle states localized within any one of these three classes of regions. Thus, the coherent propagation of wavelike quasiparticle excitations across the entire system ceases to be possible when $\omega>\Delta_{K}$, and therefore QPI is no longer present. Details of this argument, making use of a model of inhomogeneous $d$-wave superconductivity with weak scatterers present, will be presented in a future paper [69].
@@
While compelling, a crucial weakness of this picture is that it does not explain why $\Delta_{K}$ (i.e., the energy where both low-energy homogeneity and QPI disappear) exactly coincides with the energy where the tips of the contours of constant energy (CCE) of the Bogoliubov quasiparticles crosses the antiferromagnetic zone boundary. Within our model of inhomogeneous $d$-wave superconductivity, there is no reason why $\Delta_{K}$ should match this energy, which would be set by the band-structure parameters and $\overline{\Delta}$. This additional requirement likely signals that additional ingredients beyond an inhomogeneous $d$-wave order parameter should be present; this is likely where the broken-symmetry states at high energies play a key role. Other proposals for resolving the problem of QPI extinction have been put forth, invoking coexisting antiferromagnetic order [70] and modulated hoppings near impurities [71]; these proposals explain why QPI disappears when the tips of the CCEs extend past the antiferromagnetic zone boundary, but not why this energy coincides with the loss of homogeneity. A full explanation of this phenomenon will likely have to involve a combination of inhomogeneity and some of these other ingredients, in addition to squaring fully with the highly correlated nature of the underdoped superconducting state.